answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of having 3 girls?

33% for every girl... (quick but wrong)

2. for a given solo pregnancy, we have the b-g option, or 1/2.

For two pregnancies, we have bb, bg, gb, gg.

for three pregnancies, we have bbb, bbg, bgb, bgg, ggg, gbg, ggb, gbb.

Of which some outcomes are equivalent in terms of numbers. (e.g. ggb, gbg, bgg)

However only one of the eight outcomes gives three girls, so the answer is the probablilty is 1/8.

An easier method, and necessary when the number are greater, is to say the probability is 1/2 in one pregnancy.

And since (for our purposes) the pregnancies are independent events, the cumulative probability is (1/2)3 which = 1/8 or 12.5% or 0.125.

What is the theoretical probability of not rolling a 2 while the cube rolls 50 times?

The theoretical probability of not rolling a 2 while the cube rolls 50 times (calling it

event E) is: P(E) = (5/6)50 = 1.09884819... x 10-4 = 0.000109884810... ≈ 0.011%

When the value of Z -0.84 what is the area probability to the right?

-0.84= 0.2005

0.84=0.7096

so since the question is asking to the right the answer should be 0.7096

The usual sampling distributionof the difference between means is a what?

There is no such thing as "the usual sampling distribution". Different distributions of the original random variables will give different distributions for the difference between their means.

There is no such thing as "the usual sampling distribution". Different distributions of the original random variables will give different distributions for the difference between their means.

There is no such thing as "the usual sampling distribution". Different distributions of the original random variables will give different distributions for the difference between their means.

There is no such thing as "the usual sampling distribution". Different distributions of the original random variables will give different distributions for the difference between their means.

What are the chances of becoming a PE teacher?

In this economy, no job is easy to get. However, PE teachers do not require too much prior knowledge, so the odds compared to other teaching jobs are very good.

When a situation is assessed it is?

When you assess the situation, you review the problem and look over possible outcomes for the situation.

Can saliva sample testing help determine ovulation?

At-home kits that use saliva instead of urine to help determine ovulation have made it more convenient for women trying to conceive children to track their hormonal cycles, eliminating a lot of guesswork.

What are the chances of a coin landing on heads 35 times in a row?

In 34 or fewer tosses, the answer is 0. In infinitely many tosses, the answer is 1. The answer depends on the number of tosses and, since you have chosen not to share that critical bit of information, i is not possible to give a more useful answer.

Could any coin with a head on each side have been made at a mint or are they ALL novelty coins?

No, the mint has two dies to stamp the coins. One die for heads ond for tails. They fit into different parts of the machine. If you have a two headed coin you have a novelty coin. While there are no known 2-headed U.S. coins, there are 4 certified examples of genuine 2-tailed U.S. coins. And there are about 25 known genuine 2-headed or 2-tailed coins from other countries. Details about this can be seen here : http://varietynickels.com/articles/twotailed.htm The U.S. Mint designs their dies with different slots in the heads-side dies than those in the tails-side dies to prevent this error from happening. So you can see the likelihood of you having a genuine 2-headed or 2-tailed coin is infinitely small -- even less than your chances of winning the lottery. Most coins with two heads were struck at a mint the regular way, but they were cut in half and the two heads were put together by a skilled craftsman.

What is the joint probability of rolling a four on a six-sided die and correctly quessing a random picked number as being odd or even?

The two events are independent, so the answer is simply the multiple of the probabilities of the two events. So joint prob = 1/6 * 1/2 = 1/12

Probability distribution function of Hotelling's T-square distribution?

there is no pdf in hottling t sq test there is only mdf or it has multivariate distribution function

In six decks of playing cards what is the probability of six seven and eight of the same suit to come up?

The answer depends on how many cards are drawn, whether or not at random, from ordinary decks of cards, with or without replacement.

How many times would you expect to roll a number less than 5 when you toss a die 360 times?

Since a cube has 6 sides, you have a 4 in 6 chance of getting a number less than 5 each time you roll the dice. You would take how many times you want to roll it (360), divide it by the number of sides (6), and multiply it by 4 since you know you have a 4 in 6 chance. So it should be 240. This is strictly mathematical though and you should not consider it accurate since there are many variables when rolling it.

Trending Questions
How many elements are in the event space if you roll a sum of 11 with 2 dice? What is the probability of getting an ace in a pack of 52 cards? Is each element abbreviated by one or two letter? What are the odds against guessing a multiple choice answer with four choices? What is the probability of getting a red queen and a 4 on the die? How many seven digit telephone numbers are possible if the first number is not allowed to be 0 or 1? How large does a population sample need to be? What are the odds of having kids with the same birthday same time? Is it true that the probability of an event is the number of ways it can occur divided by the number of times it cannot occur? What is the frequency for heads in a series of 100 tosses of a fair coin? When are two events considered to be independent? Which shows the probability of choosing 2 blue socks from a drawer having 7 blue and 3 white socks? How can you examine events after you cleared them from event viewer? How many red and black cards are in a deck of 52 cards? What type of crash minimizes the severity of injuries through a busy intersection? You randomly select one card from a 52 card deck find the probability of selecting a nine or a two? Your friend hides a dollar bill under one of three plastic cups. Under the other two cups is a paper clip. You get to keep whatever is under the cup you choose, but you don't know which contains the dollar bill. Should you switch to the other cup? What is the theoretical probability if you hit 20 of 32 balls at bat? 46370 55170 53480 49126 89212 75292 67291 88241 37808 38154 what is the probability that a group of 5 random digits will contain at least 2 even digits? What is the entire group that the researcher is studying?